# 3D Uncertain Implicit Surface Mapping using GMM and GP

### v1

@@ IV-A Gaussian Processes
In a standard stationary GP, the prior mean function is often set to a constant number, e.g., zero, and the kernel function depends only on the distance between points in the input space and not on the specific locations of those points, e.g., the squared exponential kernel or Matérn kernel. For instance, a $3/2$ Matérn kernel used as the prior GP kernel function reads:
@@
As a stationary kernel, the hyper-parameters $\sigma_{p}$ and $l$ in the above equation should remain invariant at different input locations. In this letter, however, the prior mean and prior variance come from the outcomes of HGMM, which vary across different inputs. It indicates that the stationary GP assumption does not hold anymore and the GP regression transforms into a non-stationary case.
@@ IV-B GP Regression with Non-stationary Priors from HGMM
Figure 3: An example of selected training points for GP: (a) shows 16 Gaussian components of GMM. (b) presents the differences between estimated signed distances and measured surface points. The green color indicates small errors as shown in the color bar. The points with large errors are selected for GP training in (c). (d) denotes the GP inference results.
@@
To fully exploit the results obtained from HGMMs and reduce the model size of GP training, we select only the measurement points that are not well-modeled on the surface to be the training points. The selection criteria is given by:
@@
where $d_{m}$ is the allowed maximal discrepancy between the real observation and the HGMM estimation of surface distance. As measurements hit the surface, the real observation is always zero. This criterion indicates the HGMM estimation should be smaller than a given threshold. Otherwise, the point will be selected for GP training.
@@ IV-C GP with Derivative Observations
Therefore, we augment the original 4D training dataset with additional normal information about each point, yielding a new dataset $\mathcal{D}=\{\bm{x_{i}},y_{i},\nabla\bm{y_{i}}\}_{i=1}^{N}$. The inference can be made based on the joint Gaussian distribution of the signed distances and the gradients of distances. Accordingly, the joint covariance matrix $\tilde{k}(\bm{x},\bm{x}^{\prime})$ involves $N(6+1)$ entries[16, 29]:
@@
The gradients or the normal vectors are estimated by PCA for each point, within a small neighboring range. Given the training data $\mathcal{D}$ with derivative information, we can change the standard inference, substituting $k_{0}(\cdot,\cdot)$ in Equation (12) and (13) by Equation (17):
$\mu(\bm{x}_{*})=\begin{bmatrix}m(\bm{x}_{*})\\ \bf{0}\end{bmatrix}+{\bf\tilde{k}}_{*}^{T}({\bf\tilde{K}}_{N}+\sigma_{\eta}^{2}{\bf I})^{-1}\begin{bmatrix}\bm{y}-m({\bf X})\\ \nabla\bm{y}\end{bmatrix},$ (18)
@@
Note that, this technique has been employed in the existing work of GPIS and Log-GPIS [4, 15, 17]. They applied additional 2D observation GP and occupancy tests to estimate the normals of measurement points. Unlike that, normal vectors are estimated by PCA directly here.
@@ V Experiments and Evaluation
In this section, the datasets and experimental results are illustrated. The signed distance fields and Euclidean distance fields estimated from the proposed approach are compared with the existing state-of-the-art methods: GPIS [4] and Log-GPIS [15]. We also compare the differences between using GMM alone and GP-inference with GMM priors (GPGMM). The work is implemented in C++ and Python.
@@ V-A Dataset
The experimental evaluation of the proposed method utilizes real 3D LiDAR point clouds representing intricate urban environments, sourced from the LUCOOP dataset [30]. These point clouds were acquired through the deployment of the Riegl VMX-250 mobile mapping system, operating within Hannover, Germany. To ensure the generation of densely mapped point clouds suitable for both modeling and testing purposes, the equipped van traversed the designated route on multiple measurement campaigns. Thus, the point clouds possess sufficient density to facilitate the generation of ground truth for the signed distance field. This involves computing distances between a query point and its nearest neighbor within the reference point clouds. Approximately 10% of the entire dataset constitutes the subset employed for training purposes.
@@ V-B Practical Implementation
In the practical implementation of HGMMs, the original observations are augmented with virtual points from both sides of the surfaces. We sub-sample the point clouds and the virtual points are generated from these sub-sampled points, which is around $20\%$ of the original size. The interval between the virtual points and the hit points is set to 0.2m.
As introduced in Section IV, a 3/2 Matérn kernel is selected as the prior kernel function for GPs. The length scale parameter $l=0.08$ is set in the implementation. Large length scale equips the GP model with the capability of predicting the points at further distances, and the inference is smoother. However, the approximation can also be averaged and blurred. The smaller parameter $l$ can produce a better approximation for the points close to the surface but is only able to infer a smaller range from the surface. The determination of $l$ is a trade-off of the above considerations.
Also, the GP inference with derivatives faces scalability challenges with large datasets due to its computational complexity, which scales cubically $\mathcal{O}(N^{3}D*3)$ with the size of the data. To address this, a pragmatic implementation involves partitioning the data into blocks organized in an Octree [1] structure, the same as applied in the GPIS map [4].
@@ V-C Surface reconstruction
In Figure 5, the qualitative outcomes of the mapping process are presented alongside their corresponding uncertainties. For visualization purposes, the implicit surface function undergoes post-processing via the marching cubes algorithm[31] to render the explicit surfaces at the zero crossing of the signed distance fields. Notably, the surfaces derived from the integrated GPGMM approach (Figure 4(b)) exhibit enhanced detail compared to those generated using HGMM alone (Figure 4(a)). Furthermore, regions exhibiting a deficiency in detail correspondingly display a higher uncertainty on the GMM surface. Following the application of the GP correction, the uncertainty measures of those areas have been improved. Therefore, the results obtained from the GPGMM method show a broader distribution of areas with lower uncertainty, indicating an improvement in the overall reliability of the measurements.
Figure 5: Qualitative results extracted by the marching cubes algorithm: (a) surface of GMM; (b) surface of GPGMM. (c) and (d) are the associated uncertainty (variance) of the GMM/GPGMM surfaces; As shown by the color bars, redder colors indicate larger uncertainties and blue colors denote small uncertainties.
@@ V-D Evaluation
To evaluate both the mapping accuracy and the reliability of the uncertainty measures, the proposed GPGMM method is benchmarked with two existing state-of-the-art mapping approaches GPIS map [4] and Log-GPIS [15], which both equip uncertainty measures with the mapping results. We compared the signed distance fields of the testing points between GPIS and the proposed GMM and GPGMM. As Log-GPIS is only capable of providing Euclidean distance fields, the accuracy of EDF is compared.
There are two metrics used for evaluation: 1) Root Mean Squared Error (RMSE) for the accuracy of the distance field; 2) Log-likelihood for the uncertainty evaluation. RMSE is computed between the ground truth distances and the resultant distances from methods across a set of testing points. The log-likelihood metric quantifies how well the estimated distribution parameterized by the uncertainty measure accords with true error distribution. The larger measure indicates a better agreement between the modeled uncertainty and the true uncertainty. In this work, log-likelihood values are averaged by the number of testing samples, given by:
$LL=\frac{1}{N_{*}}\log p(f(\bf{X_{*}})),$ (20)
where $N_{*}$ is the number of the testing points $\bf{X_{*}}$.
Figure 6 and 7 provide a detailed comparative evaluation of RMSEs and log-likelihoods related to the estimated surface distances, encompassing both established benchmark methodologies and the innovative GPGMM as well as standalone GMM mapping. Both metrics are graphically represented across a spectrum of distances between the query point and the original training point clouds. With a growing distance from the observations, a decline in performance is noted across all methods, characterized by escalating RMSEs and diminishing log-likelihoods.
@@
In the analysis of the SDF accuracy among GPIS, GMM, and GPGMM (Figure 5(a)), GPGMM exhibits the best performance with the smallest RMSEs across the entire range of distances. The discrepancy between GPGMM and GMM is particularly pronounced in regions proximal to the surface. Here, GPGMM demonstrates an enhanced ability to accurately infer the surface and its adjacent distance fields - a distinction also perceptibly highlighted in Figure 5. While GPIS shows accurate inference for query points near the surface, its RMSEs notably increase as the distance from the surface increases. Mirroring this pattern, GPGMM tends to align with the GMM’s outcomes for query points located at greater distances, indicating a convergence towards the prior model in areas far from the observational data.
@@
Within the context of EDF outcomes, the performance of Log-GPIS is notably inferior, whereas GPIS generally achieves higher accuracy. For distances less than 0.1m, the outcomes from the GPGMM are marginally superior or on par with those from GPIS, but display diminished accuracy at proximities close to 0.2m. This result arises because GPIS tends to converge towards the prior beyond a certain threshold, with the experimental setup defining this prior distance as 0.2 meters. Given that signs are disregarded in EDF calculations, this prior setting intrinsically aids GPIS in obtaining accurate estimations for query points around the 0.2-meter mark. However, as distance increases, the error associated with GPIS exceeds that of the proposed GPGMM method, highlighting the limitations of relying on the fixed prior in distance estimation tasks.
@@
In the assessment of uncertainty, the proposed GPGMM demonstrates promising performance. When evaluating SDF, both the GPGMM and GMM approaches outperform GPIS, exhibiting superior likelihoods. For EDF, the uncertainty quantification provided by Log-GPIS proves to be more dependable at proximal distances, with log-likelihoods diminishing rapidly as distance increases. Overall, the GPGMM method exhibits the most consistent and robust performance relative to the alternative approaches considered.
V-E Comprehensive comparison
@@
The GPIS benchmark adopts a predetermined value for the prior mean (0.2m) and assigns the prior variance a constant value of 1. This configuration introduces limitations in the accuracy of surface estimations and uncertainty quantification. As shown in the evaluation results, it yields reasonable estimations proximate to the observations but the accuracy diminishes significantly when it predicts points located at further distances. To solve this problem, log-GPIS introduced a heat model in the standard GPIS, which modifies the regression target as the exponential transformation of the original surface distances. This converts the target value at the surface from zero into one, and transitions the prior mean from a fixed value to infinity $\sim-\log(0)$. However, Log-GPIS encounters constraints related to the choice of the length scale parameter; e.g., a small length scale parameter leads to more accurate results at the expense of surface smoothness. Additionally, the small length scale parameter also limited its capability to infer further distance, where the GP-regressed heat value $v$ reverts to the zero prior, and consequently, $d=-(l/\sqrt{3})\log(v)$ escalate rapidly to infinity outside a limited inference domain. Furthermore, the fixed variance of both GPIS and Log-GPIS results in the systematic overestimation of uncertainties.
@@
In contrast, our proposed method navigates these challenges by integrating the GMM to obtain a calibrated prior variance and distance mean, and employing local GPISs with non-stationary mean and kernels. This enhances the accuracy of signed distance field estimations and achieves more dependable uncertainty measures.
V-F Computational time comparison
@@
Regarding computational time, the proposed GPGMM and GMM reach a generally faster performance, as shown in Figure 8. Thanks to the GMM prior, the GP model size is reduced, which leads to a fast GP training and prediction process. During GP training, the time complexity of GPIS and Log-GPIS increases significantly with the number of data points, while the proposed methods remain at a low level. During the prediction phase, the performance disparities among the methods become less pronounced. Nonetheless, both GPIS and LogGPIS occasionally exhibit significantly higher computational costs, while the proposed methodologies maintain a consistently lower and more stable level of time consumption. The GMM method is often faster than GPIS and LogGPIS, suggesting its viability as a standalone approach for more time-efficient modeling, depending on the specific requirements of the applications. It is also observed that the total time complexity of GPGMM is greatly affected by the Gaussian mixture regression; A more optimized spatial structure for GMR computation will further reduce the computational time.
@@
Figure 8: Computational time: note that the orange triangle shows the GP training with only selected training points while the green triangle denotes the total computation time as the sum of GMM prior estimation and GP inference.
@@ VI Conclusions
In conclusion, this study has successfully merged the HGMM prior with GP inference to effectively model uncertain implicit surfaces in 3D spaces. This approach has extended the capabilities of previous models focused on simple building representations to encompass more complex structures and surfaces. To harmoniously integrate two probabilistic methodologies, the GMM provides efficient initial surface modeling and furnishes calibrated priors for GP inference. This integration results in a model that is not only compact but also enables the GPGMM to attain more precise estimations of surface distances and to generate more reliable uncertainty metrics due to the refined priors. Furthermore, GPGMM exhibits enhanced computational speed compared to existing GPIS and Log-GPIS frameworks, particularly noticeable during the model’s training phase.
Looking ahead, there is potential for further refinement of GMR computations through the adoption of more advanced spatial structures. This could allow for the selective use of highly accurate estimations in downstream applications, thereby optimizing performance and utility. Additionally, the methodology’s utility in localization tasks presents a promising avenue for exploration. The ability to directly leverage queried surface distances for evaluating the alignment quality between two-point clouds could significantly enhance registration processes.

### v4

@@ IV-A Gaussian Processes
In a standard stationary GP, the prior mean function is often set to a constant number, e.g., zero, and the kernel function depends only on the distance between points in the input space and not on the specific locations of those points, e.g., the squared exponential kernel or Matérn kernel. For instance, a $3/2$ Matérn kernel used as the prior GP kernel function reads as:
@@
As a stationary kernel, the hyper-parameters $\sigma_{p}$ and $l$ in the above equation should remain invariant at different input locations. In this paper, however, the prior mean and prior variance come from the outcomes of HGMM, which vary across different inputs. It indicates that the stationary GP assumption does not hold anymore and the GP regression transforms into a non-stationary case.
@@ IV-B GP Regression with Non-stationary Priors from HGMM
Figure 3: Selected training points for GP: (a) shows 16 Gaussian components of GMM. (b) presents the differences between estimated signed distances and measured surface points. The green color indicates small errors as shown in the color bar. The points with large errors are selected for GP training in (c). (d) shows the GP inference results.
@@
To fully exploit the HGMM results obtained from HGMMs and reduce the model size of GP training, we select only the points that are not well-modelled on the surface to be the training points, as shown in Figure 3. The selection criteria is given by:
@@
where $d_{m}$ is the allowed maximal discrepancy between the real observation and the HGMM estimation of surface distance. As measurements hit the surface, the real observation is always zero. This criterion indicates the HGMM estimation should be smaller than a given threshold. Otherwise, the point will be selected for GP training. This value is dependent on the required accuracy for the surface.
@@ IV-C GP with Derivative Observations
Therefore, we augment the original 4D training dataset with additional normal information about each point, yielding a new dataset $\mathcal{D}=\{\bm{x_{i}},y_{i},\nabla\bm{y_{i}}\}_{i=1}^{N}$. The inference can be made based on the joint Gaussian distribution of the signed distances and the gradients of distances. Accordingly, the joint covariance matrix $\tilde{k}(\bm{x},\bm{x}^{\prime})$ involves $N(3+1)$ entries[8, 32]:
@@
The gradients or the normal vectors are estimated by PCA for each point. Given the training data $\mathcal{D}$ with derivative information, we can change the standard inference, substituting $k_{0}(\cdot,\cdot)$ in Equation (12) and (13) by Equation (17):
$\mu(\bm{x}_{*})=\begin{bmatrix}m(\bm{x}_{*})\\ \nabla m(\bm{x}_{*})\end{bmatrix}+{\bf\tilde{k}}_{*}^{T}({\bf\tilde{K}}_{N}+\sigma_{\eta}^{2}{\bf I})^{-1}\begin{bmatrix}\bm{y}-m({\bf X})\\ \nabla\bm{y}-\nabla m(\bf{X})\end{bmatrix},$ (18)
@@
The technique of GP inference with derivative observations has been employed in the existing work [4, 8]. They applied additional 2D observation GP and occupancy tests to estimate the normals of measurement points. Unlike that, normal vectors are estimated by PCA directly here.
@@ V Experiments and Evaluation
In this section, the datasets and experimental results are illustrated. The SDF and EDF estimated from the proposed approach are compared with the existing state-of-the-art methods: GPIS [4] and Log-GPIS [8]. We also compare the differences between using GMM alone and GP-inference with GMM priors (GMMGP). The work is implemented in C++ and Python on a CPU of Intel i9-11900 @ 2.50GHz.
@@ V-A Dataset
The experimental evaluation of the proposed method utilizes real 3D LiDAR point clouds representing intricate urban environments, sourced from the LUCOOP dataset [33]. These point clouds were acquired by the Riegl VMX-250 mobile mapping system, operating within Hannover, Germany. To ensure the generation of densely mapped point clouds suitable for both modelling and testing purposes, the equipped van traversed the designated route on multiple measurement campaigns. Thus, the point clouds have a sufficient density (mean nearest distance = 5mm) to facilitate the generation of ground truth for the SDF. This involves computing distances between a query point and its nearest neighbor within the dense reference point clouds. Since the reference points are not continuous surfaces but rather discrete samples, it is essential to make them densely packed to minimize errors in the ground truth. Approximately 10% of the entire dataset is selected as training data.
@@ V-B Practical Implementation
In the practical implementation of HGMMs, the original observations are augmented with virtual points from both sides of the surfaces. To speed up training, we sub-sample the point clouds and virtual points are generated from these sub-sampled points, which is around $15\%$ of the original size. We trained the model with the subsampling rates of $5\%,15\%,40\%,80\%$, and the results show no significant differences. The HGMM approach is robust for the sparse data.
The spacing between virtual points and hit points is set to 0.25m, based on multiple considerations. The confidence in the virtual distance decreases with increasing separation from the hit point, suggesting that the designated space should not be too large. However, the small spacing restricts the model’s ability to capture distance field gradients at greater distances. Empirical results indicate that a reduced interval enhances the accuracy near the surface while leading to worse estimations at further distances. The determination of the spacing is a trade-off of the above considerations.
Also, the GP inference with derivatives faces scalability challenges with large datasets due to its computational complexity, which scales cubically $\mathcal{O}(N^{3}D^{3})$ with the size of the data. To address this, a pragmatic implementation involves partitioning the data into blocks organized in an Octree [1] structure, the same as applied in the GPIS map [4].
@@ V-C Surface reconstruction
In Figure 5, the qualitative outcomes of the mapping are presented with their corresponding uncertainties. For visualization purposes, the implicit surface undergoes post-processing via the marching cubes algorithm[34] to render the explicit surfaces at the zero crossing of the signed distance fields. Notably, the surfaces derived from the integrated GMMGP approach (Figure 5(b)) exhibit enhanced detail compared to those generated using HGMM alone (Figure 5(a)). Furthermore, regions exhibiting a deficiency in detail correspondingly display a higher uncertainty on the HGMM surface. Following the GP correction, the uncertainty measures in those areas decreased. Therefore, the GMMGP results show broader areas with lower uncertainty, indicating an improvement in the overall reliability of the model. The prior work (GMMGP2.5D) [30] uses segmented local 2.5D frames, resulting in increased surface discontinuity in reconstruction. It performs better in primarily planar areas but occasionally presents discontinuous reconstruction errors. Moreover, it generally underestimates the uncertainty relative to HGMM and GMMGP.
Figure 5: Qualitative results extracted by the marching cubes algorithm: (a) surface of HGMM; (b) surface of GMMGP; (c) surface of the previous GMMGP2.5D. (d), (e) and (f) are the associated uncertainty (variance) of the HGMM/GMMGP/GMMGP2.5D surfaces; As shown by the color bars, redder colors indicate larger uncertainties and blue colors denote small uncertainties.
@@ V-D Evaluation
To evaluate both the mapping accuracy and the reliability of the uncertainty measures, the proposed GMMGP method is benchmarked with existing state-of-the-art mapping approaches: GMMGP2.5D [30], GPIS map [4], Log-GPIS [8] and Kernel-Inverse GP (KIGP) [9], which all provide uncertainty measures along with the mapping results. The uncertainty measure of KIGP is a a proxy and does not align with our quantitative evaluation, hence it is excluded from comparison. We compared the SDF of the testing points between GMMGP2.5D, GPIS and the proposed HGMM and GMMGP. Given that Log-GPIS and KIGP only provide unsigned distance fields, EDF was compared.
There are two metrics used for evaluation: 1) Root Mean Squared Error (RMSE) for the accuracy of the distance field; 2) Predictive log-likelihood for the uncertainty evaluation. RMSE is computed between the ground truth and the resultant distances from the methods across a set of testing points. Predictive log-likelihood is a popular metric to evaluate uncertainty quality, quantifying how well the estimated distribution accords with the true distribution. The larger measure indicates a better uncertainty quality. In this work, log-likelihood values are averaged by the number of testing samples, given by:
$LL=\frac{1}{N_{*}}\sum_{i=1}^{N_{*}}\log p(\mu_{i}|\hat{\mu_{i}},\hat{\sigma}^{2}_{i})$ (20)
where $N_{*}$ is the number of the testing points $\bf{X_{*}}$, $\mu$ is the ground truth distance, $\hat{\mu}$ and $\hat{\sigma}^{2}$ are the estimation and posterior uncertainty obtained from Equation (18) and (19) .
Figure 6 and 7 present a comprehensive comparison of RMSEs and log-likelihoods for estimated surface distances, encompassing both established benchmark methodologies and our proposed GMMGP as well as standalone HGMM mapping. Both metrics are plotted against varying distances between the query points and the original training points. With a growing distance from the observations, a decline in performance is noted across all methods, characterized by increasing RMSEs and diminishing log-likelihoods.
@@
In the analysis of the SDF accuracy (Figure 6(a)), GMMGP exhibits the best performance with the smallest RMSEs across the entire range of distances. The discrepancy between GMMGP and HGMM is particularly pronounced in regions proximal to the surface. Here, GMMGP demonstrates an enhanced ability to accurately infer the surface and its adjacent distance fields - a distinction in the qualitative results is compared in Figure 5, where Figure 5(b) shows more details on the reconstructed surface than 5(a). Since the GMMGP2.5D is proposed for structured building models, more arbitrary and complicated shapes can cause improper segmentation for local planes. It shows worse robustness and accuracy than the proposed GMMGP in full 3D. Although GPIS shows accurate inference for query points near the surface, its RMSEs notably increase as the distance from the surface increases. Mirroring this pattern, GMMGP tends to align with the GMM’s outcomes for query points located at larger distances, indicating a convergence towards the prior model in areas far from the observational data.
@@
Within the context of EDF outcomes, GMMGP generally achieves higher accuracy. For distances less than 0.1m, the outcomes from the KIGP are marginally superior. At proximities close to 0.2m, GPIS shows the lowest RMSEs. This result arises because GPIS tends to converge towards the prior beyond a certain threshold, with the experimental setup defining this prior distance as 0.2 meters. Given that signs are disregarded in EDF calculations, this prior setting intrinsically aids GPIS in obtaining accurate estimations for query points around the 0.2-meter mark. However, as distance increases, the error associated with GPIS exceeds that of the proposed GMMGP method, highlighting the limitations of relying on the fixed prior in distance estimation tasks.
@@
In the assessment of uncertainty, the proposed GMMGP demonstrates promising performance. When evaluating SDF, both the GMMGP and HGMM outperform GPIS and the previous work GMMGP2.5D, exhibiting higher likelihoods. For EDF, the uncertainty quantification provided by Log-GPIS proves to be more dependable at proximal distances, with log-likelihoods diminishing rapidly as distance increases. Overall, the GMMGP method exhibits the most consistent and robust performance relative to the alternative approaches considered.
@@
The predetermined priors in GPIS lead to diminished accuracy at greater distances from observed data. To solve this problem, log-GPIS introduced a heat model in the standard GPIS, which modifies the regression target as the exponential transformation of the original surface distances. This converts the target value at the surface from zero into one, and transitions the prior mean from a fixed value to infinity $\sim-\log(0)$. However, Log-GPIS encounters constraints related to the choice of the length scale parameter; e.g., a small length scale parameter leads to more accurate results at the expense of surface smoothness. Additionally, the small length scale parameter also limited its capability to infer further distance. Although KIGP mapping employs kernel reverting functions to improve Log-GPIS, it is not scalable to the large 3D dataset as it applies full GP inference. Thus, in this case, the data must be partitioned into cells for large urban scenes, which, however, affects its accuracy.
@@
In contrast, our proposed method addresses these challenges by integrating the HGMM to obtain a calibrated prior, and employing local GPISs with non-stationary mean and kernels. This enhances the accuracy of signed distance field estimations and achieves more dependable uncertainty measures.
V-E Computation time comparison
@@
Regarding computation time, the proposed GMMGP and HGMM generally reach a faster performance, as shown in Figure 8. The introduction of the HGMM prior reduces the GP model size in GMMGP, facilitating rapid GP training nearly coinciding with HGMM. Similarly, GMMGP2.5D scales well with large datasets. However, the hierarchical structure and avoidance of segmentation in the proposed GMMGP enhance its efficiency compared to the earlier 2.5D version. During training, the time complexity of GPIS, Log-GPIS and KIGP increases significantly with the number of data points, while the proposed methods remain at a low level. This increase for GPIS and Log-GPIS is primarily attributed to their inherent normal estimation. During the prediction, KIGP incurs high computational costs, whereas performance differences among other methods are less pronounced. Nonetheless, both GPIS and LogGPIS occasionally exhibit significantly higher computational costs, while the proposed methodology maintains a consistently lower and more stable level of time consumption. The HGMM method is often faster, suggesting its viability as a standalone approach for more time-efficient modelling. It is also observed that the total time complexity of GMMGP is greatly affected by the HGMM regression; a more optimized spatial structure for GMR computation will further reduce the computational time.
@@
Figure 8: Computational time: note that blue triangles show the HGMM prior estimation while the orange triangles denote the total computation time as the sum of HGMM and GP.
@@ VI Conclusions
In conclusion, this study has successfully merged the HGMM prior with GP inference to effectively model uncertain implicit surfaces in 3D space. This approach has extended the capabilities of previous models focusing on simple building representations, to encompass more complex structures and surfaces. To harmoniously integrate two probabilistic methodologies, the GMM provides efficient initial surface modelling and furnishes calibrated priors for GP inference. This integration results in a model that is not only compact but also enables the GMMGP to attain more precise estimations of surface distances and generate more reliable uncertainty metrics due to the refined priors. Furthermore, GMMGP exhibits faster computational speed compared to other GP-based frameworks, particularly noticeable during the model’s training phase.
Looking ahead, there is potential for further refinement of GMR computations through the adoption of spatial index structures. Additionally, the method could also allow for the selective use of only accurate estimations in localization and navigation, thereby optimizing performance and utility. The methodology’s utility in localization is promising, as the queried surface distances can be directly leveraged for evaluating the alignment quality between two-point clouds in registration.